OpenStudy (anonymous):

A 70-kg diver jumps off a 10.2-m tower. (a) Find the diver's velocity when he hits the water. ?m/s (b) The diver comes to a stop 2.0 m below the surface. Find the net force exerted by the water. ?N

OpenStudy (anonymous):

Let's start with (a). We can look at solving using conservation of energy, or kinematics. Do you have an approach you want to use?

OpenStudy (anonymous):

Kinematics

OpenStudy (anonymous):

Ok. So we know vertically this will be a free-fall problem. Which gives us the known acceleration due to gravity a = 9.8 m/s^2. I will use the coordinate system where up is positive and down is negative so the acceleration would be - 9.8 m/s^2. The problem statement doesn't mention jumping upward so we will assume their initial vertical speed is 0 m/s. We also know they are coming off the 10.2m tower, so their vertical displacement will be 10.2m downward, or -10.2m viy = 0 m/s ay = -9.8 m/s^2 \[\Delta y = -10.2m\] vfy = ? So we need a kinematics equation with initial velocity, final velocity, acceleration, and displacement. \[v_{fy}^{2}=v_{iy}^{2} + 2a \Delta y\]

OpenStudy (anonymous):

How's it going?

OpenStudy (anonymous):

I'm confused by the 2aΔy part of the equation

OpenStudy (anonymous):

Final Velocity = 0 + 2(-9.8) (-10.2) Would this be right?

OpenStudy (anonymous):

yes, but remember that is the final velocity squared. so to solve you should take the square root of both sides.

OpenStudy (anonymous):

because you are taking the square root, you then should go back and check which answer would make sense a + final velocity or a - final velocity. The - makes sense with our coordinate system.

OpenStudy (anonymous):

Okay thank you! But how would you do part B?

OpenStudy (anonymous):

For part (b) you can use that final velocity as it hits the water, as your initial velocity for part be, with the diver then coming to rest (velocity final = 0 m/s) after moving 2 meters downward (-2m). You would use the same kinematics equation but this time solve for the acceleration a. Once you have the acceleration you would use Newton's 2nd law to find Force: \[\sum F = ma\]

OpenStudy (anonymous):

Which will end you with \[\Sigma F = 3498.943 N\] Thank you so much!